20. At a tea-dance, there are 4 more girls than boys. For the first dance, only couples are allowed to take the floor, i.e. girls must dance with boys and girls are not allowed to dance together. If 5/7 of the boys and 2/3 of the girls take the floor for the first dance, how many boys are there altogether?
Answer by scott8148(6628)

"there are 4 more girls than boys" ___ g = b + 4

"5/7 of the boys and 2/3 of the girls take the floor for the first dance" ___ 5b / 7 = 2g / 3 ___ 15b = 14g

substituting ___ 15b = 14(b + 4) ___ 15b = 14b + 56 ___ b = 56